How much grams of water required to evaporate from 150cm3 syrup (density of syrup = 1,258g/cm3 and procentive concentration 55%) to get the procentive concentration of sucrose 65%?
1
Expert's answer
2020-09-29T06:12:42-0400

Mass of a solution 1 = density*volume = 150cm3*1.258g/cm3 = 188.7 g

Mass of sucrose at solution 1 = 0.55*188.7 = 103.8 g

Mass of water at solution 1 = 188.7 - 103.8 = 84.9 g

Mass the solution 2 = 103.8/0.65 = 159.7 g

Mass of water at solution 2 = 159.7 - 103.8 = 55.9 g

Mass of water required to evaporate = 84.9 - 55.9 = 29 g